OpenStudy (anonymous):

*******************HELP I WILL FAN AND MEDAL*************************** A certain taxicab costs $2 per mile plus a $5 flat fee. The cab driver tells you owe him $11. How many miles m did you ride in the cab? Identify the correct equation to represent this problem and the correct solution.

OpenStudy (anonymous):

\[A. \frac{ m }{ 2 } + 5 = 11, m =12miles \] \[B. \frac{ m }{ 2 } -5=11, m = 32 miles \] \[C. 2m +5=11,m = 3 miles \] \[D. 2m - 5 =11,m = 8 miles \]

OpenStudy (anonymous):

i would say its d

OpenStudy (anonymous):

C Because 2*m+ 5 is our equation. When M = 3 miles, then 2*3 + 5 = 11 Medal?